I am confused....
Simplify. 3/4
_______
4/5

My choices: A) 3/5 B) 5/3 C) 15/16 D)16/15
I do not get any of these for my answer.

Answer by mathchemprofessor(65)

(3/4)/(4/5)
Multiplying both numerator and denominator by 5/4, the denominator equals 1.
(3/4)*(5/4)/(4/5)*(5/4)=(3/4)*(5/4)=15/16 Answer(C)